The future value of an anjuity at the end of ten years is 80000, the payments are made at the end of each year, and the interest rate is 11% compounded semiannually. What isbthe size of each annual payment?
1
Expert's answer
2021-09-07T10:16:08-0400

Here,

Given Information

Future Value of annuity at end of 10 years =80,000

Interest rate =11% compounded semi-annually

Time Period =10 years

Using the Effective Annual rate formula

"EAR=(1+\frac{r}{m})^m-1"

Where, EAR=Effective Annual rate

r=Annual nominal rate of interest

m=Number of compounding periods in a year

"EAR=(1+\frac{0.11}

{2}



)

^2

−1

\\EAR=(1+0.055)

^2

−1

\\EAR=(1.055)

^2

−1\\EAR=

1.113025−1\\

EAR=0.113025 \space or \space 11.3025\%"


Now,

Effective Annual Rate =11.3025%

Using the Future Value of annuity formula

Future Value of Annuity "=P[\frac{(1+r)

^n

−1}

r



]"

Where P=Periodic Payment 

r=rate per period 

n=number of periods 

"80000=P[\frac{(1+0.113025)^

{10}

−1}{

0.113025}\\80000\times 0.113025=





P[(1.113025)

^{10}

−1]\\

9042=P[2.9177574906−1]\\

9042=P[1.9177574906]\\

\frac{9042}{1.9177574906}=P\\

4,714.8818577530 \space or\space 4,714.88 =P"


The Size of each annual payment is 4,714.88